OpenStudy (anonymous):

PLease Help will fan and give medal! :3 Expand and then simplify the expression (a + 2)(a + 2)

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\] always put \(a=a,b=2\)

OpenStudy (anonymous):

there is no b @satellite73

OpenStudy (anonymous):

Could you show me the steps on how to solve it please?

OpenStudy (phi):

here is one way. rename the first (a+2) as A so the problem is A(a+2) now distribute the A. Can you do that ?

OpenStudy (phi):

distribute the A means multiply each term in the parens by A

OpenStudy (anonymous):

yeah I can

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

Would it be a^2+2a?

OpenStudy (phi):

I should have used a different letter than A. let's use d= (a+2) so the problem is d(a+2) now what do you get?

OpenStudy (phi):

you should get ad + 2d

OpenStudy (anonymous):

Yeah that's what I got

OpenStudy (phi):

now we remember that d is short for (a+2) and write ad+2d as a(a+2) + 2(a+2) now we have two problems that you know how to do. distribute the a and distribute the 2 all together what do you get ?

OpenStudy (anonymous):

Thank you. I'll watch it. I can understand things better if i can hear and see what's going on. Not so good online.

OpenStudy (phi):

ok, post your answer when you get it.

OpenStudy (anonymous):

a^2+4a+4

OpenStudy (phi):

yes. notice satellite's post gave the "short way" also, if we finished a(a+2) + 2(a+2) we get a^2 + 2a + 2a+ 4 and that simplifies to a^2 +4a + 4 (but if you watched the video, you know that!)

OpenStudy (anonymous):

Lol yeah I did. Thank You so much. :)